OpenStudy (anonymous):

WILL AWARD MEDAL How do you solve the inequality (x^2)(e^x) ln x > 0?? @nincompoop @Nnesha @zepdrix @jagr2713

OpenStudy (zzr0ck3r):

well \(e^x>0\) for all \(x\), \(x^2>0\) for all \(x\ne 0\) and \(\ln(x)>0\) for all \(x>1\). So ?

OpenStudy (zzr0ck3r):

hint: the \(ln(x)\) factor trumps them all.

OpenStudy (anonymous):

sorry i don't really understand where this is going.

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

These are things we can do without affecting the direction of the inequality: Add (or subtract) a number from both sides. Multiply (or divide) both sides by a positive number. Simplify a side. Hoped this helped :)

OpenStudy (zzr0ck3r):

What I am saying is that \(x^2e^x\ln(x)>0 \iff \ln(x)>0\iff x>1\)